OpenStudy (anonymous):

Help Please! A population of horses has individuals that are either gray or chestnut colored. The gray color allele, G, is dominant over the chestnut color allele, g. If the frequency of the gray allele in the population is 0.8, what percent of horses are gray? A.) 64 percent B.) 96 percent C.) 32 percent D.) 80 percent

OpenStudy (blues):

You should use the Castle Hardy Weinburg equation for this: \[1 = p^2 + 2pq + q^2\] p^2 = frequency of homozgyous dominant individuals 2pq = frequency of heterozygotes q^2 = frequency of homozgous recessive individuals. It gives you the frequency of the grey allele, p, so all you have to do is square it.